OpenStudy (anonymous):

Which trigonometric value is equal to cos 62°? A) cos 28° B) sec 28° C) sin 28° D) tan 28°

OpenStudy (hba):

By the definition of complementary trig functions: \[Sin(90-\theta)=\theta\]

OpenStudy (hba):

i mean Cos theeta*

OpenStudy (sirm3d):

uhm, @hba means \[\sin (90 - \theta)=\cos \theta\]

OpenStudy (hba):

My bad :(

OpenStudy (anonymous):

Ok thank you

OpenStudy (hba):

Your'e welcome :)